# Regular Kerr black holes: The fellowship of the string

### v1

@@ IV.3.5 The case $\epsilon_{i}=+1$ and $\epsilon_{e}=+1$
Finally, for the choice $\epsilon_{i}=\epsilon_{e}=+1$, the linear energy density and the pressure of fluid are given by $\sigma=0$ and $p=-m_{0}a/4\pi$. The linear energy density is identically zero, while the pressure can be positive or negative depending on the sign of $m_{0}$.
In this case, the roles of $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$ corresponds to the expected notion of interior and exterior regions. More importantly, let us notice that this choice for $\epsilon_{i}$ and $\epsilon_{e}$ implies that the normal points in the direction of increasing radial coordinate as seen from both regions. Since $n_{\mu}$ points from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$, the role of the radial coordinates $\rho$ and $\varrho$ stays the same with the domains given by $\rho\in[0,a]$ and $\varrho\in[a,a+\varepsilon)$ with $\varepsilon>0$, thus respecting the original domain of the radial coordinates. Therefore, this is the choice that corresponds to a consistent solution for the thin shell at $\mathcal{S}$, since it leads to a local geometry that is appropriate to match the global properties of the Kerr geometry with a regular core. Moreover, the usual assumption in order to construct regular black hole solutions is $m_{0}>0$, from which we notice that the strong energy condition is violated at the ring, what is in agreement with previous results [21].
@@ V.1 The disk and the ring
The interior spacetime $(\mathcal{M}_{i},g_{i})$ is the disk $\mathcal{D}$ with the metric given by Eq. (8) as described in Sec. II.2. Hence, the spacelike unit normal vector to $\mathcal{S}$ as seen from $\mathcal{M}_{i}$ is the same as in the preceding case, given by (15), and, moreover, the extrinsic curvature of $\mathcal{S}$ as seen from $\mathcal{M}_{e}$ is also the same as in Eq. (16).
@@ V.2 The equatorial plane and the exterior metric
Let us remind that, as the ring is approached from the outside in the equatorial plane, the metric given by Eq. (11) reduces to the form (12). In contrast with the metric given by Eq. (13) used by the authors in Ref. [7], here we consider the full metric that comes from using $m(\varrho)\approx m_{0}(\varrho^{2}-a^{2})^{3/2}$ for $\varrho\to a$ in Eq. (11), without any further approximation. In other words, besides taking $\theta=\pi/2$ into the full metric, the only approximation we make is by choosing the asymptotic form of the mass function $m(\varrho)$.
Therefore, the exterior spacetime $(\mathcal{M}_{e},\,g_{e})$ now refers to the equatorial region outside the ring $\mathcal{A}$ with the metric given by Eq. (12). The spacelike unit normal vector to $\mathcal{S}$ as seen from $\mathcal{M}_{e}$ is given by (17), and, moreover, the extrinsic curvature of $\mathcal{S}$ is given by
$K^{t}_{et}=-2m_{0}a\epsilon_{e},\ \ K^{t}_{e\phi}=2m_{0}a^{2}\epsilon_{e},$
$K^{\varphi}_{et}=-2m_{0}\epsilon_{e},\ \ K^{\varphi}_{e\varphi}=\frac{\epsilon_{e}}{a}+2m_{0}a\epsilon_{e}.$ (20)
@@ V.3.1 The matter-energy content at the ring
Thus, we can see that, differently from the EMT given by Eq. (19), in this case the EMT of the shell is not diagonal. In fact, the presence of nondiagonal terms like $\mathcal{S}^{t}_{\varphi}$ is an indication that the fluid at the shell is rotating with respect to the asymptotic observer.
V.3.2 Orientations of the normal vector in which $\epsilon_{i}=-\epsilon_{e}$
If $\epsilon_{i}=-\epsilon_{e}$, we can diagonalize the EMT of the shell such that the eigenvalues are given by
@@ V.3.2 Orientations of the normal vector in which $\epsilon_{i}=-\epsilon_{e}$
with the respectively normalized eigenvectors $v^{a}_{\pm}$ given by
$v^{t}_{\pm}=\sqrt{a^{2}\left(v^{\varphi}_{\pm}\right)^{2}\pm 1},$ (23)
$v^{\varphi}_{\pm}=\left[\pm\big{(}\omega^{-2}_{\pm}-a^{2}\big{)}\right]^{-1/2},$ (24)
@@
$\omega_{\pm}$ $=\frac{S^{\varphi}_{\ t}}{\lambda_{\pm}-S^{\varphi}_{\ \varphi}}=\frac{2m_{0}a}{1+2m_{0}a^{2}\pm\sqrt{1+4m_{0}a^{2}}}.$ (25)
The timelike eigenvector that corresponds to the comoving observer with the shell is $v^{a}_{+}$, while the spacelike eigenvector is $v^{a}_{-}$. Hence, we can define the energy density of the shell as $\sigma=-\lambda_{+}=-S_{ab}v^{a}_{+}v^{b}_{+}$, while $p=\lambda_{-}=S_{ab}v^{a}_{-}v^{b}_{-}$ is the tangential pressure. Moreover, the eigenvector $v^{a}_{+}$ can be decomposed as
$v^{a}_{+}=\gamma\left(\delta^{a}_{\ t}+\omega_{+}\delta^{a}_{\ \varphi}\right),$ (26)
where $\gamma=\omega_{+}^{-1}/(\omega^{-2}_{+}-a^{2})^{1/2}$ and $\omega_{+}$ can be interpreted as the angular velocity of the rotating shell.
@@
For $\epsilon_{e}=-1$, which also means $\epsilon_{i}=+1$, the energy density of the shell is positive, while the pressure can be either positive or negative depending on the sign of $m_{0}$. Similarly to the case discussed in Sec. IV.3.2, this scenario represents a closed universe with a rotating string at the ring joining the two regions.
For $\epsilon_{e}=1$, which also means $\epsilon_{i}=-1$, the energy density of the shell is negative, while the pressure can be either positive or negative depending on the sign of $m_{0}$. Moreover, similar to the case discussed in Sec. IV.3.3, this scenario corresponds to an open universe with a rotating string at the ring joining the two regions.
V.3.3 Orientations of the normal vector in which $\epsilon_{i}=\epsilon_{e}$
@@
If $\epsilon_{i}=\epsilon_{e}$, it is not possible to diagonalize the EMT of the shell since the determinant of $S^{a}_{\ b}$ vanishes and, therefore, we are in the presence of a null dust fluid. This is indeed the case once that we can recast $S^{a}_{\ b}$ as
@@
In fact, in the present case it is a constant, and can be interpreted as the energy density of the shell. Hence, the thin-shell ring corresponds to a rotating string composed by a null dust fluid.
In the case with $\epsilon_{e}=1$, which also means $\epsilon_{i}=1$, the energy density depends on the sign of $m_{0}$, being positive for $m_{0}<0$. Notice that a negative $m_{0}$ means a negative mass function. Moreover, $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$ correspond to the expected notion of interior and exterior regions, respectively, with both regions being joined by a rotating string at the ring. Hence, this is the appropriate choice to match the overall geometry of the Kerr-like spacetime with a regular rotating core. However, there is a drawback regarding the kind of matter of the string, the energy density of the string is negative for $m_{0}>0$. It is worth remembering that, when the extension through the region $r<0$ is not performed, the metric components of the Kerr spacetime are not smooth in the region $r=0$, what Israel [30] interpreted as caused by a layer of mass spread over the disk. The material distributed over the disk presents negative energy density and rotates with superluminal velocity. Here, the situation is not so extreme since the string is composed by a null dust with negative energy density. This implies that the null energy condition is violated at the ring.
In the case with $\epsilon_{e}=-1$, which also means $\epsilon_{i}=-1$, the energy density depends on the sign of $m_{0}$, being positive for $m_{0}>0$. Similarly to the case discussed in Sec. IV.3.4, the roles of $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$ are swapped and the resulting geometry does not correspond to the original problem.
@@ VI Discussion
Following the proposal by Smailagic and Spallucci [7], in which the problematic central ring of the Kerr-like geometries is replaced with a string, we analyzed all possible thin shell configurations that can be built from the Darmois-Israel formalism regarding the possible orientations of the normal vector relative to the radial coordinate. The different orientations of the normal vector engender four different configurations and may be parameterized by two unit parameters $\epsilon_{e}=\pm 1$ and $\epsilon_{e}=\pm 1$.
Considering the original choice $\epsilon_{i}=-\epsilon_{e}=+1$ made by Smailagic and Spallucci [7], we conclude that the resulting solution is not appropriate for a thin shell at the ring of a rotating regular black hole, once the resulting spacetime topology contradicts the original domain of the exterior radial coordinate and the notion of the exterior region as the equatorial region just outside the ring. In fact, such a choice gives rise to a closed spacetime, that may be interpreted as a Minkowski-de Sitter closed universe, corresponding to a bubble universe.
@@
More importantly, we formally demonstrated that the only consistent solution for the thin shell at the ring of a rotating regular black hole results from the choice $\epsilon_{i}=\epsilon_{e}=+1$. In this case, the original domains of the radial coordinates are respected and the roles of the interior and exterior regions are preserved. Hence, differently from Ref. [7], we find that the suitable string solution replacing the Kerr ring has a vanishing linear energy density and a pressure that can be either positive or negative depending on the sign of the constant $m_{0}$.
@@
Finally, we noticed that in the scenario considered in Ref. [7], the string of matter replacing the ring is not rotating. In order to remedy this fact, we constructed a scenario where the string is indeed rotating, and considered all the four combinations of the normal vector orientations. The results for the corresponding geometries are the same as for the nonrotating string, while the kinds of material result quite different. In the cases where $\epsilon_{i}=-\epsilon_{e}$, the rotating string corresponds to a fluid with positive energy density for $\epsilon_{i}=-\epsilon_{e}=-1$, and with negative energy density for $\epsilon_{i}=-\epsilon_{e}=+1$. In the case where $\epsilon_{i}=\epsilon_{e}$, the rotating string corresponds to a null fluid with positive energy density for $m_{0}\epsilon_{e}<0$, and with negative energy density for $m_{0}\epsilon_{e}>0$.

### v2

@@ IV.3.5 The case $\epsilon_{i}=+1$ and $\epsilon_{e}=+1$
Finally, for the choice $\epsilon_{i}=\epsilon_{e}=+1$, the line energy density and the pressure of the fluid are given by $\sigma=0$ and $p=-m_{0}a/4\pi$. The line energy density is identically zero, while the pressure can be positive or negative depending on the sign of $m_{0}$.
In this case, the features of $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$ correspond to the expected notion of the interior and exterior regions. More importantly, let us notice that this choice for $\epsilon_{i}$ and $\epsilon_{e}$ implies that the normal points in the direction of increasing radial coordinates, as seen from both regions. Since $n_{\mu}$ points from $\mathcal{M}_{i}$ to $\mathcal{M}_{e}$, the role of the radial coordinates $\rho$ and $\varrho$ stays the same with the domains given by $\rho\in[0,a]$ and $\varrho\in[a,a+\varepsilon)$ with $\varepsilon>0$, thus respecting the original domain of the radial coordinates. Therefore, this is the choice that corresponds to a consistent solution for the thin shell at $\mathcal{S}$, since it leads to a local geometry that is appropriate to match the global properties of the Kerr geometry with a Minkowski disk glued to a regular dS(AdS) region outside the ring, as illustrated in Fig. 4. Moreover, the usual assumption in order to construct regular black hole solutions is $m_{0}>0$, from which we notice that the strong energy condition is violated at the ring, which is in agreement with previous results [22].
Figure 4: Embedding diagram of a $t=\text{constant}$ and $\theta=\pi/2$ slice of $(\mathcal{M}_{i},g_{i})$ with $\epsilon_{i}=+1$ and $(\mathcal{M}_{e},g_{e})$ with $\epsilon_{e}=+1$ in $3$-dimensional Euclidean space. The domain of the interior radius is $0\leq\rho\leq a$. The domain of the exterior radius is $a\leq\varrho<\infty$. The string is represented by the black solid lines in each spacetime, with the lines being identified through the junction.
@@ V.1 The disk and the ring
[…]
@@ V.2 The equatorial plane and the exterior metric
Let us recall that, as the ring is approached from the outside in the equatorial plane, the metric given by Eq. (22) reduces to the form (24). In contrast with the metric given by Eq. (25) used by the authors in Ref. [8], here we consider the full metric that comes from using $m(\varrho)=m_{0}(\varrho^{2}-a^{2})^{3/2}$ for $\varrho\to a$ in Eq. (22), without any further approximation. In other words, in addition to taking $\theta=\pi/2$ into the full metric, the only approximation we make is choosing the asymptotic form of the mass function $m(\varrho)$.
[…]
$K^{t}_{(e)t}=-2m_{0}a\epsilon_{e},\ \ K^{t}_{(e)\varphi}=2m_{0}a^{2}\epsilon_{e},$
$K^{\varphi}_{(e)t}=-2m_{0}\epsilon_{e},\ \ K^{\varphi}_{(e)\varphi}=\frac{\epsilon_{e}}{a}+2m_{0}a\epsilon_{e}.$ (34)
We now have all the ingredients to perform the matching between the two spacetime regions.
@@ V.3.1 The matter-energy content at the ring
Thus, we can see that, different from the EMT given by Eqs. (31) and (32), in this case the EMT of the shell is not diagonal. In fact, the presence of nondiagonal terms such as $\mathcal{S}^{t}_{\ \varphi}$ is an indication that the fluid at the shell is rotating with respect to the asymptotic observer.
As we have seen in the previous section, the appropriate interpretation of the energy density and pressure, as well as the resulting topology of the complete spacetime, depends on the particular values of $\epsilon_{i}$ and $\epsilon_{e}$, and then it is interesting to study each one of the fourth combinations separately.
V.3.2 The cases in which $\epsilon_{i}=-\epsilon_{e}$
We first study the interpretation of the EMT (35)by considering $\epsilon_{i}=-\epsilon_{e}$.
The characteristic equation for the matrix (35) give us $\lambda^{2}-2\epsilon\lambda/a-4m_{0}=0,$ whose solutions are the eigenvalues
@@ V.3.2 The cases in which $\epsilon_{i}=-\epsilon_{e}$
and the respective normalized eigenvectors are
$\mathfrak{v}_{\pm}^{a}=N_{\pm}\left(\frac{1+2m_{0}a^{2}\pm\sqrt{1+4m_{0}a^{2}}}{2m_{0}\,a}\delta_{t}^{a}+\delta_{\varphi}^{a}\right),$ (37)
@@
$N_{\pm}^{-2}=\pm\left[a^{2}-\left(\frac{1+2m_{0}a^{2}\pm\sqrt{1+4m_{0}a^{2}}}{2m_{0}a}\right)^{2}\right].$ (38)
Considering that $m_{0}$ may assume negative values, we find three different situations for the eigenvalues $L_{\pm}$ depending on the discriminant $D=1+4m_{0}a^{2}$: Case (i) $D>0$, case (ii) $D=0$, and case (iii) $D<0$. We analyze each case separately.
Case (i): If $01+4m_{0}a^{2}>$ 0, the two eigenvalues $\lambda_{\pm}$ assume real values and the canonical form of the EMT (35) is
$S_{\ a}^{b}=\sigma u_{a}u^{b}+px_{a}x^{b}=\left(\sigma+p\right)u_{a}u^{b}+ph_{a}^{b},$ (39)
where
$\begin{split}&\sigma=-\lambda_{+},\qquad u^{a}=v_{+}^{a}\\ &p=\lambda_{-},\;\;\ \qquad x^{a}=v_{-}^{a},\end{split}$ (40)
with $u^{a}$ being a timelike vector and $x^{a}$ being a spacelike vector.
Moreover, the timelike eigenvector $u^{a}=v^{a}_{+}$ can be decomposed in terms of the coordinate basis as
$u^{a}=\gamma\left(\delta^{a}_{\ t}+\omega\delta^{a}_{\ \varphi}\right),$ (41)
where $\gamma=\left(1-\omega^{2}a^{2}\right)^{-1/2}$ and
$\omega$ $=\frac{2m_{0}a}{1+2m_{0}a^{2}+\sqrt{1+4m_{0}a^{2}}}.$ (42)
can be interpreted as the angular velocity of the matter fluid along the rotating string.
The stress-energy tensor (39) can be interpreted as a perfect fluid. For $\epsilon_{e}=-1$, which also means $\epsilon_{i}=+1$, the energy density of the shell $\sigma$ is positive, while the pressure $p$ is negative. For $\epsilon_{e}=+1$, which also means $\epsilon_{i}=-1$, the energy density of the shell $\sigma$ is negative, while the pressure $p$ is positive.
Case (ii): If $01+4m_{0}a^{2}=$ 0, the two eigenvalues $\lambda_{\pm}$ are degenerated (identical) and the canonical form of the EMT (35) is
$S_{\ a}^{b}=\sigma\left(u_{a}u^{b}+x_{a}x^{b}\right)+\beta k_{a}k^{b},$ (43)
where
$\begin{split}&\sigma=-p=-\frac{\epsilon_{e}}{8\pi a},\qquad\beta=-\frac{\epsilon_{e}}{16\pi a},\\ &k^{a}=\delta_{\ t}^{a}-\frac{1}{a}\delta_{\ \varphi}^{a},\ \qquad u^{a}=\delta_{\ t}^{a},\qquad x^{a}=\frac{1}{a}\delta_{\ \varphi}^{a},\end{split}$ (44)
with $k^{a}$ being a lightlike vector, $u^{a}$ being a timelike vector, and $x^{a}$ being a spacelike vector.
The stress-energy tensor (43) can be interpreted as a mixture of a perfect fluid with a flux of directed radiation (lightlike fluid) along the string. For $\epsilon_{e}=-1$, which also means $\epsilon_{i}=+1$, the energy density of the matter fluid $\sigma$ and the energy density of the lightlike fluid $\beta$ are both positive, while the pressure $p=-\sigma$ of the matter fluid is negative. For $\epsilon_{e}=+1$, which also means $\epsilon_{i}=-1$, the energy density of the matter fluid $\sigma$ and the energy density of the lightlike fluid $\beta$ are both negative, while the pressure $p=-\sigma$ of the matter fluid is positive.
Case (iii): If $01+4m_{0}a^{2}<$ 0, the two eigenvalues $\lambda_{\pm}$ assume complex values, and the canonical form of the EMT (35) is
$S^{a}_{\ b}=\sigma\left(u^{a}u_{b}-x^{a}x_{b}\right)+\beta\left(u^{a}x_{b}+x^{a}u_{b}\right),$ (45)
where
$\begin{split}&\sigma=-\frac{\epsilon_{e}}{8\pi\,a},\qquad p=-\sigma=\frac{\epsilon_{e}}{8\pi\,a},\\ &\beta=-\frac{\epsilon_{e}}{8\pi\,a}\sqrt{|1+4m_{0}a^{2}|},\\ &u^{a}=\sqrt{\frac{1}{2}+\Big{|}\frac{m_{0}a}{8\pi\beta}\Big{|}}\,\delta_{\ t}^{a}+\frac{1}{a}\sqrt{\Big{|}\frac{m_{0}a}{8\pi\beta}\Big{|}-\frac{1}{2}}\,\delta_{\ \varphi}^{a}\\ &x^{a}=\sqrt{\Big{|}\frac{m_{0}a}{8\pi\beta}\Big{|}-\frac{1}{2}}\,\delta_{\ t}^{a}+\frac{1}{a}\sqrt{\frac{1}{2}+\Big{|}\frac{m_{0}a}{8\pi\beta}\Big{|}}\,\delta_{\ \varphi}^{a},\end{split}$ (46)
with $u^{a}$ being a unit timelike vector and $x^{a}$ being a unit spacelike vector.
The stress-energy tensor (45) can be interpreted as a (perfect) fluid with heat flow along the string, with $u^{a}$ representing the four-velocity of the fluid and $x^{a}$ giving the direction of the heat flow. For $\epsilon_{e}=-1$, which also means $\epsilon_{i}=+1$, the energy density of the shell $\sigma$ is positive, while the pressure $p$ is negative. Since $\beta>0$, the heat flux is along $x^{a}$ spacelike direction. For $\epsilon_{e}=+1$, which also means $\epsilon_{i}=-1$, the energy density of the shell $\sigma$ is negative, while the pressure $p$ is positive. The change from $\beta>0$ to $\beta<0$ indicates the reversion the heat flux with respect to $x^{a}$.
@@
Now we consider the topology of the resulting spacetime in the two cases engendered by the choice $\epsilon_{i}=-\epsilon_{e}$. For $\epsilon_{e}=-1$, which also means $\epsilon_{i}=+1$, the situation is identical to case discussed in Sec. IV.3.2, whose resulting topology is depicted in Fig. 1. This scenario represents a closed universe with a rotating string at the ring joining the two regions. For $\epsilon_{e}=-1$, which also means $\epsilon_{i}=+1$, the situation is identical to case discussed in Sec. IV.3.3, whose resulting topology is depicted in Fig. 2. This scenario corresponds to an open universe with a rotating string at the ring joining the two regions.
V.3.3 The cases in which $\epsilon_{i}=\epsilon_{e}$
@@
If $\epsilon_{i}=\epsilon_{e}$, it is not possible to diagonalize the EMT of the shell since the determinant of $S^{a}_{\ b}$ vanishes and, therefore, the EMT can be interpreted as a lightlike dust fluid. This is indeed the case once that $S^{a}_{\ b}$ may be recast as
@@
In fact, in the present case, $H$ is a constant and can be interpreted as the energy density of the shell. Hence, the thin-shell ring corresponds to a string composed of a rotating lightlike dust that rotates with angular velocity $1/a$.
In the case with $\epsilon_{e}=-1$, which also means $\epsilon_{i}=-1$, the energy density of the string depends on the sign of $m_{0}$, which is positive for $m_{0}>0$. Similarly to the case discussed in Sec. IV.3.4, the roles of $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$ are swapped, and the resulting geometry does not correspond to the original problem. See also Fig. 3.
In the case with $\epsilon_{e}=1$, which also means $\epsilon_{i}=1$, the energy density of the string depends on the sign of $m_{0}$, which is positive for $m_{0}<0$. Moreover, $\mathcal{M}_{i}$ and $\mathcal{M}_{e}$ correspond to the expected notion of interior and exterior regions, respectively, with both regions being joined by a rotating string at the ring (see also Fig. 4). Hence, this is an appropriate choice to match the overall geometry of the Kerr-like spacetime with a regular rotating core. However, there is a drawback regarding the kind of matter of the string. The energy density of the string is negative for $m_{0}>0$, while the center core of the Kerr-like geometry presents a rotating de Sitter-type fluid with a positive energy density. Interestingly, this case is similar to the string source investigated by Israel in Ref. [36] for the Kerr geometry,in which the ring is composed of a rotating dust-like material that also rotates with angular velocity $1/a$. In such a work by Israel, the energy density is negative in the region $r<0$. Here, we do not investigate the extension through the region $r<0$ for the Kerr-like geometry. On the other hand, the choice $m_{0}<0$ leads to a string with a positive energy density and the center core of the Kerr-like geometry contains a rotating anti-de Sitter-type fluid.
@@ VI Discussion
We studied the central region of regular Kerr-like geometries with a variable mass function that depends on the radial Boyer-Lindquist radius as $m(r)\simeq r^{3}$. First, we reviewed the behavior of the main curvature scalars close to $r=0$ and confirmed that such scalars present a jump at the Kerr ring. Then, following the proposal by Smailagic and Spallucci [8], in which the problematic central ring in the Kerr-like geometries is replaced with a string, whose interior region is flat and whose exterior region is described by a de Sitter-type metric, we studied two different cases.
@@
The first case analyzed is the same spacetime construction studied in [8], but here considering all possible thin shell configurations that can be built from the Darmois-Israel formalism, regarding the possible orientations of the normal vector relative to the radial coordinate. The different orientations of the normal vector engender four different configurations and can be parameterized by two unit parameters $\epsilon_{e}=\pm 1$ and $\epsilon_{e}=\pm 1$.
@@
Considering the original choice $\epsilon_{i}=-\epsilon_{e}=+1$ made by Smailagic and Spallucci [8], we conclude that the resulting solution is not appropriate for a thin shell at the ring of a rotating regular black hole, once the resulting spacetime topology contradicts the original domain of the exterior radial coordinate and the notion of the exterior region as the equatorial region just outside the ring. In fact, such a choice gives rise to a closed spacetime that may be interpreted as a Minkowski-de Sitter closed universe, corresponding to a bubble universe. More importantly, we formally demonstrated that the only consistent solution for the thin shell at the ring of a rotating regular black hole results from the choice $\epsilon_{i}=\epsilon_{e}=+1$. In this case, the original domains of the radial coordinates are respected and the roles of the interior and exterior regions are preserved. Hence, differently from Ref. [8], we find that the suitable string solution replacing the Kerr ring has a vanishing line energy density and a pressure that can be either positive or negative depending on the sign of the constant $m_{0}$.
Finally, we noticed that in the scenario considered in Ref. [8], the string of matter replacing the ring is not rotating. This result is obtained by verifying that both the angular velocity of the matter and the frame-dragging effect in the string are vanishing.
@@
The second case analyzed here is a spacetime construction similar to that of the first case, but now the exterior metric is different from that considered in [8]. More specifically, we constructed a new scenario in which the string is indeed rotating and considered the four possible combinations of normal vector orientations. The results for the corresponding geometries are the same as for the nonrotating string, while the kinds of material result quite different. In cases where $\epsilon_{i}=-\epsilon_{e}$, the matter content of the rotating string corresponds to a fluid with a positive energy density for $\epsilon_{i}=-\epsilon_{e}=-1$, and with a negative energy density for $\epsilon_{i}=-\epsilon_{e}=+1$. In the case where $\epsilon_{i}=\epsilon_{e}$, the rotating string corresponds to a lightlike fluid with a positive energy density for $m_{0}\epsilon_{e}<0$, and with a negative energy density for $m_{0}\epsilon_{e}>0$. Since the combination that leads to an appropriate solution with respect to the resulting spacetime topology is $\epsilon_{e}=\epsilon_{i}=1$, we have two possible Kerr-like solutions that depends on the sign of $m_{0}$. For $m_{0}<0$, the energy density of the string is positive, which means a negative mass function for the regular Kerr geometry and is interpreted as an anti-de Sitter-type of matter. For $m_{0}>0$, the energy density of the string is negative, however the mass function of the regular Kerr geometry is positive and is interpreted as a de Sitter-type of matter.